# Gaga: Group Any Gaussians via 3D-aware Memory Bank

### v2

@@ 4.1 Experimental Setup
SA3D [3]* 24.9 23.8 7.4 7.0 42.5 39.2
LERF [14]* 33.5 30.6 28.3 14.7 49.7 42.6
LEGaussians [25] 34.6 32.6 31.4 18.8 42.8 35.3
LangSplat [24] 61.9 60.9 61.9 54.7 59.8 52.7
Gaussian Grouping [32]* 69.7 67.9 77.0 68.7 71.7 66.1
OmniSeg3D [33] 85.0 83.7 83.6 75.5 69.8 63.8
@@
Table 1: Quantitative results for 3D query tasks on LERF-Mask. Gaga outperforms previous approaches, showcasing favorable performance in mIoU and BIoU with SAM. * denotes the results are reported in [32].
@@
4.2 Open-vocabulary 3D Query on LERF-Mask
We compare our method with 6 state-of-the-art methods on 3D scene understanding: LERF [14], LEGaussians [25], LangSplat [24], SA3D [3], Gaussian Grouping [32] and OmniSeg3D [33]. The first three methods focus on CLIP feature embedding. We calculate the relevancy between rendered CLIP features and query text features following their official implementation. The last three methods leverage 2D segmentation masks predicted by SAM [16]. Since SAM does not support language prompts, SA3D, Gaussian Grouping and Gaga adopt Grounding DINO [19] to identify the mask ID in a 2D image and pick the corresponding 3D mask. OmniSeg3D does not provide mask IDs, so we use Grounding DINO to identify a pixel that semantically aligns with the text prompt and select the corresponding 3D mask based on a similarity threshold of 0.9.
Tab. 1 illustrates that Gaga achieves superior results in mIoU and mBIoU compared to previous methods utilizing SAM as 2D segmentation method. As shown in Fig. 5, the visualization results of 3D query tasks with prompts “paper napkin” and “bag of cookies” further demonstrate the advancement of Gaga, as Gaga provides more clear segmentation masks without artifacts.
@@ 4.2 Open-vocabulary 3D Query on LERF-Mask
Table 2: Quantitative results on Replica and ScanNet. Gaga performs well with both 2D segmentation methods on two datasets. The performance of Gaussian Grouping (noted as GG) varies significantly with different 2D segmentation methods, whereas Gaga consistently delivers stable performance.
Figure 6: Qualitative results on Replica and ScanNet. Gaga consistently predicts more accurate segmentation masks. Gaussian Grouping (noted as GG) often covers the same object with different masks (row 1, 3 and 5), creating large empty regions (row 1, 2 and 3), or misidentifying similar instances (row 4 and 5).
@@
Table 3: Quantitative results on Replica with limited data. Gaga consistently outperforms Gaussian Grouping with both 2D segmentation methods. The percentage of IoU drop shows that Gaga has greater robustness against reductions in training data.
4.3 3D Segmentation on Replica and ScanNet
Tab. 2 presents the quantitative comparison results on Replica and ScanNet datasets. Gaga exhibits better performance on both datasets regardless of the 2D segmentation model used, showcasing its stability across different datasets and models and consistently achieving high performance. Qualitative results are shown in Fig. 6. Row 1 and 2 depict the visualizations from the Replica dataset, while rows 3-5 showcase results from the ScanNet dataset. Gaussian Grouping [32] frequently assigns different mask IDs to the same object, resulting in inconsistent masks and empty regions. Row 4 shows that Gaussian Grouping struggles to distinguish similar objects, whereas Gaga accurately identifies each object by leveraging 3D information.
4.4 3D Segmentation with Limited Data on Replica
To demonstrate the robustness of Gaga against changes in training image quantity, we sparsely sample the Replica training set with ratios of 0.3, 0.2, 0.1, and 0.05. As depicted in Tab. 3, Gaga consistently exhibits superior performance in terms of IoU, with approximately a 15% advantage using EntitySeg and a 25% advantage using SAM. Remarkably, when utilizing SAM, Gaga surpasses fully trained Gaussian Grouping with just 5% of the training data by more than 10% (31.87% vs. 21.76%). We also compute the IoU drop compared to using all training images by:
$IoU\ Drop(x\%)=\frac{IoU(100\%)-IoU(x\%)}{IoU(100\%)},$ (6)
where $IoU(x\%)$ denotes the IoU achieved when $x\%$ of the training data is used. Compared to Gaussian Grouping, Gaga exhibits less sensitivity to decreases in the number of training images, as evidenced by smaller values in IoU drop. Visual results are shown in Fig. 7. With just 5% of the training data, Gaga delivers accurate segmentation masks, whereas Gaussian Grouping fails to provide masks for a significant portion of objects due to inaccurate tracking.
Figure 7: Qualitative results on Replica with limited data. The visualizations depict samples when using only 5% of the training data. Even with limited data, Gaga consistently produces high-quality segmentation. In contrast, Gaussian Grouping struggles to track objects accurately and leaves significant empty regions.
@@
We further showcase the performance of Gaga on the diverse MipNeRF 360 dataset. We provide visualization comparison on both indoor and outdoor scenes with Gaussian Grouping [32] in Fig. 8, using SAM [16] (row 1, 2) and EntitySeg [23] (row 3, 4). We display two images for each scene to assess the consistency across different views. Gaga offers more detailed segmentation, while segmentation masks generated by Gaussian Grouping exhibit severe artifacts and empty regions. Additionally, inconsistency across two views exists in the rendering results of Gaussian Grouping, as shown in row 1 and row 3.
@@ 4.6 Application: Scene Manipulation
Figure 9: Scene manipulation results on MipNeRF 360 and Replica. Gaga accurately identifies the cushion part of the footstool, whereas Gaussian Grouping colors it entirely. For object removal and translation tasks, Gaga generates more precise 3D entities with fewer artifacts, resulting in better visual performance.
Figure 10: Visual comparison of mask association methods. Gaga with 3D-aware memory bank achieves a superior visual quality and closer to the ground truth. Notice that the Video Tracker baseline mislabels the wall and floor, Our Memory Bank (w. All Gaussians) mislabels the floor.
Gaga achieves high-quality and multi-view consistent 3D segmentation, beneficial for tasks like scene manipulation. We demonstrate the effectiveness of this application on the MipNeRF 360 [1] and Replica [27] datasets (see Fig. 9). In the left image, we change the color of the footstool’s cushion and remove the stuffed animal on the armchair. Gaga accurately identifies 3D Gaussians representing the cushion, while Gaussian Grouping fails, coloring the entire footstool maroon along with part of the sofa and some floating Gaussians. Gaga also effectively groups and removes the entire part of the stuffed animal on the sofa with minimal artifacts, whereas Gaussian Grouping leaves many floating Gaussians. Similar results are observed in experiments involving the position shifting of a chair in the right image.
@@ 4.7 Ablation Study on Mask Association Method
SAM (Upper Bound) [16] 60.89 57.07 67.16
Linear Assignment [26] 1.92 1.54 0.58
@@
Video Tracker [32] 21.76 25.00 19.72
Our Memory Bank (All Gaussians) 42.30 40.48 46.51
@@
Table 4: Ablation study on mask association methods. Our mask association method with the 3D-aware memory bank surpasses the previous video tracker baseline on all metrics.
We conduct ablation studies to evaluate the effectiveness of the proposed mask association method on the Replica dataset. The baselines for comparison include: (1) Linear Assignment: Cost-based linear assignment mask association proposed in Panoptic Lifting [26]. (2) w/o. Mask Association: Lifting inconsistent 2D masks to 3D. (3) Video Tracker[32] is employed as a representative method. (4) Our Memory Bank (w. All Gaussians): same as Gaga except that it selects all Gaussians splatted to the mask as its corresponding Gaussians. (5) Our Memory Bank: i.e., Gaga. We also add a baseline SAM (Upper Bound), which uses SAM to process rendered RGBs from Gaussian Splatting and evaluate on each single frame without considering multi-view consistency. This baseline can serve as an upper bound to show the inherent difference between class-agnostic and panoptic segmentation.
Quantitative results in Tab. 4 indicate that Gaga with the 3D-aware memory bank achieves superior performance compared to the previous method with video tracker. The cost-based linear assignment method used in [26] is not suitable for class-agnostic segmentation tasks, as it involves a significantly larger number of masks compared to panoptic segmentation. Comparison with Our Memory Bank (w. All Gaussians) baseline demonstrates the effectiveness of using depth maps to select corresponding 3D Gaussians for each mask. We also show the visual comparison in Fig. 10.
5 Conclusions
We introduce Gaga, a framework that reconstructs and segments open-world 3D scenes using inconsistent 2D masks predicted by zero-shot segmentation models. Gaga employs a 3D-aware memory bank to categorize 3D Gaussians and establishes mask association across different views by identifying the overlap between Gaussians projected to each mask. Results on various datasets show that Gaga outperforms previous methods with superior segmentation accuracy, multi-view consistency, and reduced artifacts. Additional application in scene manipulation further highlights Gaga’s high segmentation accuracy and practical utility.
@@ 5 Conclusions
1 Overview
In this supplementary document, we provide further experimental results, including more results on scene manipulation and sparse view setting in Sec. 3. We then delve into more experimental details of the datasets, metrics and implementation in Sec. 4. More ablation studies are shown in Sec. 5 and limitations are discussed in Sec. 6.
2 Supplementary Video
Please watch the supplementary demo video for a comprehensive introduction and visual comparison between our method Gaga and the current state-of-the-art methods. The video features additional qualitative comparisons and an animation illustration of Gaga.
Figure 11: Scene manipulation results on MipNeRF 360 and LERF-Mask. Gaga accurately identifies the flowerpot without affecting the color of the plant. Notice that Gaussian Grouping [32] creates a cyan region on the wooden door behind. For the object removal and duplication tasks, Gaga can also provide more accurate results with fewer artifacts.
3 Supplementary Experimental Results
3.1 Results on Scene Manipulation
@@
We provide additional results for the downstream scene manipulation task to further demonstrate the prospect of applying Gaga to real-world scenarios. On the “counter” scene of the MipNeRF 360 dataset [1], we change the color of the flowerpot to cyan and duplicate the glass jar. Gaussian Grouping [32] can not differentiate the plant and flowerpot, whereas Gaga generates a more accurate segmentation mask. Additionally, Gaga produces a clearer boundary and avoids artifacts on the iron tray when duplicating the glass jar.
In the “figurines” scene of the LERF-Mask dataset [32], we transform the yellow duck to blue and remove the red toy chair. Gaga precisely changes only the duck’s color without affecting other objects, and achieves a more thorough removal of the red toy chair.
3.2 Results on Sparsely Sampled Replica
@@
4 Experimental Details
4.1 Details on Datasets
We employ the official script from Gaussian Splatting [13] for colmap to acquire camera poses and the initial point cloud. Consequently, the actual number of images utilized in the experiment might be lower than expected due to colmap process failures. Please refer to Tab. 5 for the scene names used in the Replica and ScanNet datasets.
LERF-Mask Dataset [32]. LERF-Mask is based on the LERF dataset [14] and annotated with tasks and ground truth by the author of [32]. It contains 3 scenes: figurines, ramen, and teatime. For each scene, 6-10 objects are selected as text queries, and Grounding DINO [19] is utilized to select the mask ID from the rendered segmentation.
Replica Dataset [27]. We select 8 scenes from the entire Replica Dataset the same as [35]. We use the rendered results provided by authors of [35] and follow their data processing process: for each scene, we uniformly select 20% images as training data and 20% images as test data from all rendered RGB images. This results in 180 training images and 180 test images for each scene.
@@ 4.1 Details on Datasets
ScanNet Dataset [10]. DM-NeRF [29] selects 8 scenes from the entire ScanNet dataset. Each scene has approximately 300 images for training and about 100 images for testing. We utilize 7 out of the 8 scenes, excluding ”scene 0024_00” due to the subpar 3D reconstruction results in both Gaussian Splatting [13] and Gaussian Grouping [32].
MipNeRF 360 Dataset [1]. We downsample the images by a factor of 4, consistent with the setting in [32], to accommodate the large size of the original images. For novel view synthesis evaluation, we set the sample step at 8, the same as the setting in [13].
@@
Table 5: Selected scenes in Replica and ScanNet. We select 8 scenes from the Replica dataset following [35], and 7 scenes from the ScanNet dataset following [29].
4.2 Details on Evaluation Metrics
@@
         iou_matrix[$i$][$j$] $\leftarrow$ iou_list.mean()
@@
   paired_iou $\leftarrow$ iou_matrix[$gt\_indices$][$pred\_indices$]
@@
4.3 Further Implementation Details
For training vanilla 3D Gaussians, we maintain the same parameter setting as [13]. To train the identity encoding, we freeze all the other attributes of Gaussians and use the same parameter setting as [32]. The identity encoding has 16 dimensions, and the rendered 2D identity encoding is in the shape of 16 $\times$ $h$$\times$ $w$, where $h$ and $w$ represent the height and width of the image. The classifier for predicting mask ID given the 2D identity encoding and selecting Gaussians for editing given the 3D identity encoding shares the same architecture, with 16 input channels. The number of output channels equals the number of groups in the 3D-aware memory bank after associating all images. All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU.

### v3

@@ 4.2 Open-vocabulary 3D Query on LERF-Mask
SA3D sa3d * 24.9 23.8 7.4 7.0 42.5 39.2
LERF lerf * 33.5 30.6 28.3 14.7 49.7 42.6
LEGaussians legaussians 34.6 32.6 31.4 18.8 42.8 35.3
LangSplat langsplat 61.9 60.9 61.9 54.7 59.8 52.7
Gaussian Grouping gg * 69.7 67.9 77.0 68.7 71.7 66.1
OmniSeg3D omniseg3d 85.0 83.7 83.6 75.5 69.8 63.8
@@
* denotes results reported in gg .
We compare our method with six state-of-the-art methods on 3D scene understanding: LERF lerf , LEGaussians legaussians , LangSplat langsplat , SA3D sa3d , Gaussian Grouping gg and OmniSeg3D omniseg3d . The first three methods focus on CLIP feature embedding. We calculate the relevancy between rendered CLIP features and query text features following their official implementation. The last three methods leverage 2D segmentation masks predicted by SAM sam . Since SAM does not support language prompts, SA3D, Gaussian Grouping and Gaga adopt Grounding DINO grounding_dino to identify the mask ID in a 2D image and pick the corresponding 3D mask. OmniSeg3D does not provide mask IDs, so we use Grounding DINO to identify a pixel that semantically aligns with the text prompt and select the corresponding 3D mask based on a similarity threshold of 0.9. Tab. 1 illustrates that Gaga achieves superior results in mIoU and mBIoU compared to previous methods utilizing SAM as 2D segmentation method. As shown in Fig. 5, the visualization results of 3D query tasks with prompts “paper napkin" and “bag of cookies" further demonstrate the advancement of Gaga, as Gaga provides more clear segmentation masks without artifacts.
@@
Figure 6: Qualitative results on Replica and ScanNet. Gaga consistently predicts more accurate segmentation masks. Gaussian Grouping often covers the same object with different masks (row 1, 3, 5), creating large empty regions (row 1, 2, 3), or misidentifying similar instances (row 4, 5).
4.3 3D Segmentation on Replica and ScanNet
Table 2: Quantitative results on Replica and ScanNet. Gaga performs well with both 2D segmentation methods on two datasets, while the performance of Gaussian Grouping varies significantly with different segmentation methods.
@@
Tab. 2 presents the quantitative comparison results on Replica and ScanNet datasets. Gaga exhibits better performance on both datasets regardless of the 2D segmentation model used, showcasing its stability across different datasets and models and consistently achieving high performance. Qualitative results are shown in Fig. 6. Row 1 and 2 depict the visualizations from the Replica dataset, while rows 3-5 showcase results from the ScanNet dataset. Gaussian Grouping gg frequently assigns different mask IDs to the same object, resulting in inconsistent masks and empty regions. Row 4 shows that Gaussian Grouping struggles to distinguish similar objects, whereas Gaga accurately identifies each object by leveraging 3D information.
Figure 7: Qualitative results on Replica with limited data. The visualizations depict samples when using only 5% of the training data, where Gaga still produces high-quality segmentation. In contrast, Gaussian Grouping struggles to track objects accurately and leaves significant empty regions.
4.4 3D Segmentation with Limited Data on Replica
Table 3: Results on Replica with limited data. Gaga outperforms Gaussian Grouping with both 2D segmentation methods. IoU drop shows that Gaga has greater robustness against reductions in training data.
@@
To demonstrate the robustness of Gaga against changes in training image quantity, we sparsely sample the Replica training set with different ratios. Depicted in Tab. 3, Gaga consistently exhibits superior performance in terms of IoU, with approximately a 15% advantage using EntitySeg and a 25% advantage using SAM. Remarkably, when utilizing SAM, Gaga surpasses fully trained Gaussian Grouping with just 5% of the training data by more than 10% (31.87% $v.s.$ 21.76%). We also compute the IoU drop compared to using all training images by: $IoU\ Drop(x\%)=\frac{IoU(100\%)-IoU(x\%)}{IoU(100\%)}.$ where $IoU(x\%)$ denotes the IoU achieved when $x\%$ of the training data is used. Gaga exhibits less sensitivity to decreases in the number of training images, as evidenced by smaller values in IoU drop. Visual results are shown in Fig. 7. With just 5% of the training data, Gaga delivers accurate segmentation masks, whereas Gaussian Grouping fails to provide masks for a significant portion of objects due to inaccurate tracking.
@@
We further showcase the performance of Gaga on the diverse MipNeRF 360 dataset. We provide visualization comparison on both indoor and outdoor scenes with Gaussian Grouping gg in Fig. 8, using SAM sam (row 1, 2) and EntitySeg entityseg (row 3, 4). We display two images for each scene to assess the consistency across different views. Notably, inconsistency across the two views exists in the rendering results of Gaussian Grouping, as shown in row 1 and 3.
@@ 4.6 Application: Scene Manipulation
Figure 9: Scene manipulation results. Gaga accurately identifies the cushion part of the footstool, whereas Gaussian Grouping colors it entirely. For object removal and translation tasks, Gaga generates more precise 3D entities with fewer artifacts, resulting in better visual performance.
Figure 10: Visual comparison of mask association methods. Gaga with 3D-aware memory bank achieves a superior visual quality and is closer to the ground truth. Notice that the Video Tracker baseline mislabels the wall and floor. Our Memory Bank (w. All GS) mislabels the floor.
Gaga achieves high-quality and multi-view consistent 3D segmentation, beneficial for scene manipulation tasks. We demonstrate the effectiveness of this application on the MipNeRF 360 mipnerf and Replica replica datasets in Fig. 9. We change the color of the footstool’s cushion and remove the stuffed animal from the armchair. Gaga accurately identifies 3D Gaussians representing the cushion, while Gaussian Grouping fails, coloring the entire footstool maroon along with part of the sofa and some floating Gaussians. Gaga also effectively groups and removes the entire part of the stuffed animal on the sofa with minimal artifacts, whereas Gaussian Grouping leaves many floating Gaussians. Similar results are observed in experiments involving the position shifting of a chair in the right image.
@@ 4.7 Ablation Study on Mask Association Method
Table 4: Ablation study. Our 3D-aware memory bank surpasses the previous video tracker baseline on all metrics.
@@
SAM (Upper Bound) sam 60.89 57.07 67.16
Linear Assignment panoptic_lifiting 1.92 1.54 0.58
@@
Video Tracker (gg, ) 21.76 25.00 19.72
Our Memory Bank (All GS) 42.30 40.48 46.51
@@
We conduct ablation studies to evaluate the effectiveness of the proposed mask association method on the Replica dataset. The baselines include: (1) Linear Assignment: Cost-based linear assignment mask association proposed in Panoptic Lifting panoptic_lifiting . (2) w/o. Mask Association: Lifting inconsistent 2D masks to 3D. (3) Video Trackergg is employed as a representative method. (4) Our Memory Bank (w. All GS): same as Gaga except that it selects all Gaussians splatted to the mask as its corresponding Gaussians. (5) Our Memory Bank: i.e., Gaga. We also add a baseline SAM (Upper Bound), which uses SAM to process rendered RGBs from Gaussian Splatting and evaluate on each single frame without considering multi-view consistency. This baseline can serve as an upper bound to show the inherent difference between class-agnostic and panoptic segmentation. Quantitative results in Tab. 4 indicate that Gaga with the 3D-aware memory bank achieves superior performance compared to the previous method with video tracker. The cost-based linear assignment method used in panoptic_lifiting is not suitable for class-agnostic segmentation tasks, as it involves a significantly larger number of masks compared to panoptic segmentation. Comparison with Our Memory Bank (w. All Gaussians) baseline demonstrates the effectiveness of using depth maps to select corresponding 3D Gaussians for each mask. Visual comparison is shown in Fig. 10.
5 Conclusions and Limitations
Gaga reconstructs and segments open-world 3D scenes using inconsistent 2D masks predicted by zero-shot segmentation models. Gaga employs a 3D-aware memory bank to categorize 3D Gaussians and establishes mask association across different views by identifying the overlap between Gaussians projected to each mask. Results on various datasets show that Gaga outperforms previous methods with superior segmentation accuracy, multi-view consistency, and reduced artifacts. Although Gaga could leverage closed-set 2D panoptic segmentation masks to enable 3D panoptic segmentation, we leave this extension to future work beyond the scope of this paper.
@@ 5 Conclusions and Limitations
Appendix A Overview
In this supplementary document, we provide further experimental results, including more results on scene manipulation and sparse view setting in Sec. B. We then delve into more experimental details of the datasets, metrics and implementation in Sec. C. More ablation studies are shown in Sec. D and limitations are discussed in Sec. E.
Figure 11: Scene manipulation results on MipNeRF 360 and LERF-Mask. Gaga accurately identifies the flowerpot without affecting the color of the plant. Notice that Gaussian Grouping (gg, ) creates a cyan region on the wooden door behind. For the object removal and duplication tasks, Gaga can also provide more accurate results with fewer artifacts.
Appendix B Supplementary Experimental Results
B.1 Results on Scene Manipulation
@@
We provide additional results for the downstream scene manipulation task to further demonstrate the prospect of applying Gaga to real-world scenarios. On the “counter" scene of the MipNeRF 360 dataset mipnerf , we change the color of the flowerpot to cyan and duplicate the glass jar. Gaussian Grouping gg can not differentiate the plant and flowerpot, whereas Gaga generates a more accurate segmentation mask. Additionally, Gaga produces a clearer boundary and avoids artifacts on the iron tray when duplicating the glass jar.
In the “figurines" scene of the LERF-Mask dataset gg , we transform the yellow duck to blue and remove the red toy chair. Gaga precisely changes only the duck’s color without affecting other objects, and achieves a more thorough removal of the red toy chair.
B.2 Results on Sparsely Sampled Replica
@@
Appendix C Experimental Details
C.1 Details on Datasets
We employ the official script from Gaussian Splatting 3dgs for colmap to acquire camera poses and the initial point cloud. Consequently, the actual number of images utilized in the experiment might be lower than expected due to colmap process failures. Please refer to Tab. 5 for the scene names used in the Replica and ScanNet datasets.
LERF-Mask Dataset gg . LERF-Mask is based on the LERF dataset lerf and annotated with tasks and ground truth by the author of gg . It contains 3 scenes: figurines, ramen, and teatime. For each scene, 6-10 objects are selected as text queries, and Grounding DINO grounding_dino is utilized to select the mask ID from the rendered segmentation.
Replica Dataset replica . We select 8 scenes from the entire Replica Dataset the same as semantic_nerf . We use the rendered results provided by authors of semantic_nerf and follow their data processing process: for each scene, we uniformly select 20% images as training data and 20% images as test data from all rendered RGB images. This results in 180 training images and 180 test images for each scene.
@@ C.1 Details on Datasets
ScanNet Dataset scannet . DM-NeRF dm_nerf selects 8 scenes from the entire ScanNet dataset. Each scene has approximately 300 images for training and about 100 images for testing. We utilize 7 out of the 8 scenes, excluding "scene 0024_00" due to the subpar 3D reconstruction results in both Gaussian Splatting 3dgs and Gaussian Grouping gg .
MipNeRF 360 Dataset mipnerf . We downsample the images by a factor of 4, consistent with the setting in gg , to accommodate the large size of the original images. For novel view synthesis evaluation, we set the sample step at 8, the same as the setting in 3dgs .
Table 5: Selected scenes in Replica and ScanNet. We select 8 scenes from the Replica dataset following (semantic_nerf, ), and 7 scenes from the ScanNet dataset following (dm_nerf, ).
@@
C.2 Details on Evaluation Metrics
@@
   $i$][$j$] $\leftarrow$ iou_list.mean()
@@
[…]
@@
C.3 Further Implementation Details
For training vanilla 3D Gaussians, we maintain the same parameter setting as 3dgs . To train the identity encoding, we freeze all the other attributes of Gaussians and use the same parameter setting as gg . The identity encoding has 16 dimensions, and the rendered 2D identity encoding is in the shape of 16 $\times$ $h$$\times$ $w$, where $h$ and $w$ represent the height and width of the image. The classifier for predicting mask ID given the 2D identity encoding and selecting Gaussians for editing given the 3D identity encoding shares the same architecture, with 16 input channels. The number of output channels equals the number of groups in the 3D-aware memory bank after associating all images. All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU.